OpenStudy (anonymous):

The perimeter or a rectangle is 10 times the width. The length is 9 metres more than the width. FInd the width of the rectangle?

OpenStudy (anonymous):

Let the width =x then length will be (x+9) perimeter =2[x+(x+9)]=4x+18 therefore we have equation 10x=4x+18 on solving it x=3